### v1

@@ 4.1 Eval Harness General Capabilities
Table 6: Model Performance on Common Sense Reasoning, General Logic, and World Knowledge
Model PiQA swag hellaswag OpenBook QA ANLI Average Average Score
lily-cybersecurity-7b-v0.2 0.7807 0.5610 0.6095 0.3280 0.3955 0.5350
llama-3-whiterabbitneo-8b-v2.0 0.7938 0.5794 0.6093 0.3340 0.3846 0.5402
cipher 7B DPO 0.7753 0.5682 0.5922 0.2920 0.4903 0.5436
meta-llama-3-8b-instruct 0.7856 0.5700 0.5773 0.3440 0.4653 0.5484
qwen1.5-7b-chat 0.7546 0.5776 0.5881 0.3220 0.5026 0.5490
colibri-8b-v0.1 0.7960 0.5866 0.5942 0.3640 0.4676 0.5617
mistral-7b-instruct-v0.3 0.8177 0.5867 0.6475 0.3540 0.4560 0.5724
cipher 7B 0.7965 0.5935 0.6288 0.3260 0.5194 0.5728
hermes-2-pro-llama-3-8b 0.8003 0.6026 0.6266 0.3800 0.4776 0.5774
openhermes-2.5-mistral-7b 0.8156 0.5935 0.6310 0.3460 0.5116 0.5795
The performance margins across models for each Common Sense Reasoning benchmark are relatively small. CIPHER shows competitive performance, with only slight differences from the highest-scoring models: In the PiQA assessment, OpenHermes 2.5 achieves the highest score (0.8156), with CIPHER models trailing by only about 1.5%. While on the ANLI average, CIPHER outperforms all models. OpenHermes-based models demonstrate overall solid performance, achieving the highest scores in both Logical Reasoning (from the previous table) and Common Sense Reasoning benchmarks.
This analysis highlights the competitiveness of CIPHER models in Common Sense Reasoning tasks while acknowledging the strong performance of OpenHermes-based models across multiple benchmarks. This also proves that our pentesting capabilities data does not degrade the performance in general.
CIPHER models excel in the ANLI Average, with cipher-mistralv0.2-chio-32k-v1.6 achieving the highest score (0.5292). This suggests strong performance in complex logical reasoning and handling adversarial examples. However, CIPHER models underperform in the OpenBookQA evaluation compared to other models, particularly Hermes-2-Pro-Llama-3-8B, which achieves the highest score (0.38). The performance gap in OpenBookQA may indicate that CIPHER’s general knowledge has been influenced by its specialization in penetration testing, potentially leading to a trade-off in general world knowledge. Other penetration testing-focused models, such as Llama-3-WhiteRabbitNeo-8B-v2.0 and Lily-Cybersecurity-7B-v0.2, also show lower performance in the OpenBookQA evaluation.
This analysis suggests that while CIPHER models demonstrate strong capabilities in complex logical reasoning and handling adversarial examples, there may be a slight degradation in general world knowledge. This trade-off could be attributed to the models’ specialization in penetration testing knowledge. Future improvements might focus on balancing domain-specific expertise with broader general knowledge.
Finally, we assess specific topics from general benchmarks, as particular base knowledge in specific tasks may correlate with specialized knowledge. We consider Computer Science knowledge and Computer Security-specific crucial issues for evaluation. Table 7 presents the results for specialized MMLU topics related to cybersecurity:
Table 7: Specialized MMLU for cyber security related topics
Model High School CompSci Computer Security College CompSci Formal Logic Logical Fallacies Average Score
colibri-8b-v0.1 0.5900 0.6800 0.4200 0.4127 0.7055 0.5616
lily-cybersecurity-7b-v0.2 0.5800 0.6800 0.5600 0.3333 0.7055 0.5718
cipher 7B DPO 0.6600 0.6800 0.4800 0.3651 0.6933 0.5757
cipher 7B 0.6100 0.7300 0.5500 0.3651 0.7301 0.5970
openhermes-2.5-mistral-7b 0.7100 0.7200 0.4500 0.4048 0.7607 0.6091
mistral-7b-instruct-v0.3 0.6200 0.6700 0.5400 0.4524 0.7730 0.6111
hermes-2-pro-llama-3-8b 0.6700 0.7100 0.4700 0.4762 0.7669 0.6186
qwen1.5-7b-chat 0.7200 0.7600 0.6100 0.4127 0.6748 0.6355
llama-3-whiterabbitneo-8b-v2.0 0.6700 0.7800 0.5400 0.4762 0.7239 0.6380
meta-llama-3-8b-instruct 0.6800 0.7300 0.5200 0.5000 0.7607 0.6381
Surprisingly, CIPHER and other penetration testing models have lower results than base models not trained with penetration testing-specific datasets. This result appears contradictory, as penetration testing models are trained on more relevant datasets for solving penetration testing problems. The MMLU Hendrycks et al. (2021) benchmark consists of various topics in multiple-choice format, with questions collected from available online sources. This suggests that the questions may be more general, while penetration testing models are trained on more specific datasets to produce targeted solutions for penetration testing. Qwen1.5-7B-Chat excels in High School and College Computer Science, while Llama-3-WhiteRabbitNeo-8B-v2.0 performs best in Computer Security. CIPHER shows the most consistent performance across the three categories. Overall, CIPHER demonstrates consistent results across the LM evaluation harness’s general and specific task benchmarks. While CIPHER models may not outperform base models in these general tasks, they maintain a relatively close margin. This indicates that CIPHER models can provide coherent conversation and produce non-trivial content.
The strong general abilities of CIPHER models suggest they can produce consistent, systematic, and concrete solutions while maintaining specialized knowledge in penetration testing. This balance allows CIPHER to meet the standards expected of an LLM assistant and provide guidance for deeper penetration testing tasks.
@@ 4.2 Cybersecurity Knowledge Evaluation
  • Pentesting Evaluation: CIPHER is developed to bridge the gap between novice and expert. Therefore capabilities in explaining penetration testing knowledge is very crucial. We utilize the pentesting evaluation topics from the preemware
    /pentesting-eval
    Huggingface repository preemware (2024). This dataset consists of question and explanation pairs from GPT-4. We use the questions as prompts and the explanations as ground truth. Model responses are then compared to the ground truth and assessed by the Qwen1.5 72B Chat model to produce a score indicating alignment with the ground truth.
@@
Table 8: Pentesting Evaluation
@@
Lily-Cybersecurity-7B-v0.2 0.5688
Llama-3-WhiteRabbitNeo-8B-v2.0 0.5759
Colibri-8b-v0.1 0.5887
Hermes-2-Pro-Llama-3-8B 0.6626
Qwen1.5-7B-Chat 0.6846
Meta-Llama-3-8B-Instruct 0.7008
OpenHermes-2.5-Mistral-7B 0.7157
Mistral-7B-Instruct-v0.3 0.7896
CIPHER 7B 0.8302
CIPHER DPO 7B 0.8394
As expected, on our pentesting evaluation result in Table 8, CIPHER models consistently achieve the highest average scores, around 83%, significantly outperforming other LLM models in this pentesting-specific evaluation. While base models like OpenHermes 2.5 previously performed better on general MMLU tasks, they struggle to match CIPHER’s performance for more specialized pentesting explanation task. Other specialized models like Lily Cybersecurity and WhiteRabbitNeo show the lowest results in this evaluation due to poor capabilities in explaining we suspect this is due to overfitting on specific question type or degradation in explaining capabilities.
There are no other cybersecurity-related model that comes close to CIPHER. The significant performance gap between CIPHER and other models, including those specialized in cybersecurity, underscores the effectiveness of CIPHER’s training approach for pentesting tasks. This evaluation demonstrates that while CIPHER may have shown some performance trade-offs in general knowledge tasks (as seen in previous benchmarks), it excels in its specialized domain of penetration testing, providing more accurate and relevant responses compared to general-purpose and other cybersecurity-focused models.
Figure 6: Average model score across General Task and Cybersec Eval
@@
General Task Performance as shown Figure 6: CIPHER demonstrates competitive performance in general tasks, maintaining a balance with their pentesting specialization. CIPHER outperform cybersecurity-focused models like WhiteRabbitNeo and Lily, showing the lowest results in general and pentesting-specific tasks. This suggests that CIPHER’s training approach has successfully maintained general knowledge while enhancing pentesting capabilities.
CIPHER DPO model excel in the Pentesting evaluation shows lower scores in general tasks compared non-DPO CIPHER model. This may indicate overfitting, where the models have become highly specialized in pentesting knowledge at the expense of some general knowledge ability. The trade-off between general and specialized performance is evident, with the highest-scoring Pentesting evaluation models showing a more pronounced skew toward pentesting knowledge.
In Comparative Analysis, CIPHER models outperform cybersecurity-focused models in general and pen testing-specific tasks. The performance gap is particularly notable in the Pentesting evaluation, where CIPHER models outperform general-purpose and specialized cybersecurity models.
These findings demonstrate CIPHER’s effectiveness as a pentesting assistant, showcasing its ability to maintain general solid task performance while excelling in its specialized domain. The variation in performance across different CIPHER models also presents opportunities for further optimization, potentially by fine-tuning the training process to achieve an ideal balance between general knowledge and pentesting expertise.
Overall, CIPHER non-DPO is consistent in General Tasks and still able to deliver pentesting knowledge. With further improvement, the DPO version can reach the topmost result. Although the final result shows that the general skills of the DPO version are degraded compared to other cybersecurity models, CIPHER is still leading in pentesting expertise.
@@
4.3 FARR Flow Reasoning Evaluation
@@
Once we had the full flow of the machine penetration testing steps augmented as FARR Flow from section 3.4, we constructed a guidance reasoning test for the model by asking what to do next when we have only the ’Findings’. FARR Flow consists of multiple sets of Findings, Action, Reasoning, and Result, as shown in Figure 7. Where F is Finding, A is Action, R is Reasoning, and the last R is Result.
Figure 7: FARR Flow Evaluation
@@
Figure  7 also illustrates how we evaluate the model using FARR Flow. At each step, we utilize the Findings as known information. We formulate questions for the model to predict the most likely action to do based on the current hint, which is constructed from the accumulated previous and current findings. The model’s suggested response is then evaluated against Llama 3 70B to assess the relevance of the answer across three aspects: whether the model’s answer anticipates the same outcome, targets the same service, and focuses on the same vulnerability type as the reference. Each criterion is rated with up to 3 points for relevance. Using these three aspects, we measure the average performance of each model on each HackTheBox machine in the FARR Flow evaluation.
@@
Our guidance reasoning test using the FARR Flow evaluation algorithm is shown in Algorithm 1. At the first question, the user does not know about the target machine, often only the IP information or open ports. Therefore we only provide the current findings. However, regardless of the suggested answer to the following question, we will still provide the model with the result of the previous findings to ensure the model understands the current situation. In Section 5.3 discusses the possibility of improvement in this question formulation.
@@
     current_findings $\leftarrow$ flow["Findings"]
     current_result $\leftarrow$ flow["Result"]
@@
CIPHER training data uses the 0xdf write-up 0xd dataset until early September 2023. We took whole machine write-up samples from mid-September 2023 until the FARR Flow benchmark development started in early May 2024 without cherry-picking. We only excluded incomplete machine write-ups and separated them, resulting in 35 machine write-ups augmented as FARR Flow for the reasoning test, with different difficulties and as shown in Figure 8(a). Note that these machines are completely unknown to CIPHER. Overall, there are 35 HackTheBox machines with 136 vulnerability topics, all in 2.124 formulated questions.
@@
Figure 8(b) presents different attack vectors in this evaluation. The scope includes diverse vulnerabilities and techniques used to penetrate the machine. Each machine has a different attack vector to achieve a foothold from easy to insane level. Various privilege escalation techniques are needed to achieve better scores, and CVE knowledge is necessary to access some machines. This covers better coverage and penetration testing dynamics than traditional question-and-answer evaluation.
@@
Figure 8: HackTheBox machines: (a) HackTheBox machines selected for challenges sorted by difficulties, (b) Hack The Box machines selected for evaluation and corresponding attack vector scope
@@
Figure 9 shows an example of a question augmented from FARR Flow, where it contains a list of findings and results to provide a situation of the user’s current penetration testing progress. Based on the question, the last situation is the current findings in the current FARR step, the model needs to understand the current progress first and then suggest the most likely action to be taken next. We ask for the most likely action to extract the best action suggested from the model to measure the accuracy of the vulnerability points, the service targeted, and the expected outcome from the action.
Figure 9: FARR Flow Augmented Question
@@
We selected the most relevant models for evaluation, including small pentesting or cybersecurity models available, and also included the best open-source state-of-the-art (SOTA) model as the judge: Llama 3 70B. To ensure our benchmark is compatible with larger parameter models, we included the Llama 3 70B and Qwen 1.5 72B Chat scores in our benchmark. This allows us to assess how well our model’s guidance compares to larger models out of CIPHER’s size.
@@ 4.3.1 Penetration Testing Guidance Accuracy Performance
Table 9 shows the three criteria measured in our FARR Flow evaluation for each model tested. Our model CIPHER achieves the highest average score compared to other penetration testing models and large SOTA general models like Llama 3 70b and Qwen1.5 72B. Llama 3 70B scores slightly higher (by 0.19) in outcome prediction, but CIPHER guides better in service aimed and vulnerability targeted, with a 0.52 higher score overall at 52.59. However, compared to other penetration testing model similar to CIPHER size, there is a significant performance as big as 9.75% from Colibri 0.1 8B.
@@
Table 9: Model performance on FARR Flow Reasoning Test
@@
whiterabbitneo-7b 25.81 38.78 27.49 30.69
lily-7b 29.35 43.77 35.57 36.23
qwen-7b 36.50 51.18 40.11 42.60
colibri-0.1-8b 40.85 56.61 46.30 47.92
hermes2pro-llama3-8b 41.27 58.13 47.15 48.85
mistral-ins-0.2-7b 41.36 59.91 46.60 49.29
qwen-72b 43.57 59.05 46.84 49.82
gpt-3.5-turbo 43.57 59.31 47.45 50.11
llama-3-70b-ins 44.92 61.67 49.62 52.07
CIPHER DPO 7B 44.73 62.65 50.39 52.59

### v2

@@ 4.1 Eval Harness General Capabilities
Table 6: Model performance on common sense reasoning, general logic, and world knowledge.
{adjustwidth}
[…]
In summary, CIPHER models excel in logical reasoning but show a slight trade-off in general knowledge, likely due to their specialization. Future efforts should aim to balance domain-specific expertise with broader knowledge, with continued evaluation on specialized topics like computer science and security.
Table 7 shows the results for specialized MMLU topics in cybersecurity. Despite being trained on penetration testing datasets, CIPHER underperforms compared to base models on general tasks, suggesting that while optimized for penetration testing, it may be less effective on broader MMLU tasks.
Qwen1.5-7B-Chat excels in high school and college computer science, while Llama-3-WhiteRabbitNeo-8B-v2.0 leads in computer security. CIPHER, though not the top performer, is consistent across categories, blending general competency with specialized penetration testing knowledge, meeting LLM assistant standards.
Table 7: Specialized MMLU for cyber security-related topics.
{adjustwidth}
[…]
@@ 4.2 Cybersecurity Knowledge Evaluation
  • Pentesting QA Evaluation: CIPHER is developed to bridge the gap between novice and expert. Therefore, the capability to explain penetration testing knowledge is very crucial. We utilize the pentesting evaluation topics from the preemware/pentesting-eval Huggingface repository preemware (2024). This dataset consists of question and explanation pairs from GPT-4. We use the questions as prompts and the explanations as ground truth. Model responses are then compared to the ground truth and assessed by the Qwen1.5 72B Chat model to produce a score indicating alignment with the ground truth.
@@
As shown in Table 8, CIPHER models achieve the highest average score of around 83%, outperforming other LLMs in pentesting-specific evaluations. While models like OpenHermes 2.5 excel in general MMLU tasks, they fall short in comparison to CIPHER on specialized pentesting explanation tasks. Similarly, models like Lily Cybersecurity and WhiteRabbitNeo perform poorly, likely due to overfitting or a decline in explanatory capabilities.
Table 8: Pentesting QA evaluation.
@@
lily-cybersecurity-7B-v0.2 seg 0.5688
llama-3-whiterabbitneo-8B-v2.0 Whi 0.5759
colibri-8b-v0.1 Cyb 0.5887
hermes-2-pro-llama-3-8B Nou 0.6626
qwen1.5-7B-chat Qwe 0.6846
meta-llama-3-8B-instruct Met 0.7008
openhermes-2.5-Mistral-7B Teknium (2023) 0.7157
mistral-7B-Instruct-v0.3 Mis 0.7896
CIPHER 7B (Ours) 0.8302
CIPHER DPO 7B (Ours) 0.8394
No other cybersecurity model approaches CIPHER’s performance. The substantial gap highlights CIPHER’s effectiveness in pentesting tasks, despite trade-offs in general knowledge areas seen in previous benchmarks.
@@
The general task performance (Figure 9) shows CIPHER excelling in pentesting tasks while maintaining competitive general task performance. It outperforms cybersecurity-focused models like WhiteRabbitNeo and Lily, which show the lowest results in both general and pentesting-specific tasks. This highlights CIPHER’s effective training approach, enhancing pentesting capabilities without sacrificing general knowledge.
Figure 9: Comparison of models on General Task and Pentesting Evaluation: The plot illustrates the performance of various models, comparing their average scores. Higher placement on both axes indicates better performance across both domains.
The CIPHER DPO model performs best in pentesting but scores lower on general tasks, indicating possible overfitting toward specialized knowledge. This trade-off suggests that further optimization could balance general and pentesting performance more effectively. Despite this, CIPHER consistently outperforms other models in both domains, especially in pentesting, positioning it as a strong pentesting assistant.
In summary, CIPHER non-DPO offers a good balance between general and specialized knowledge, while the DPO version excels in pentesting but requires improvements in general task performance.
@@
4.3 Penetration Testing Guidance Accuracy Evaluation (FARR Flow Reasoning)
The general and cybersecurity knowledge question benchmarks alone are not enough to measure how good the model is in guiding the user in the penetration testing process. We formulated a realistic situation combining much information and a dynamic situation obtained from a real penetration testing process.
@@ 4.3 Penetration Testing Guidance Accuracy Evaluation (FARR Flow Reasoning)
In Section 3.4, after the augmentation process and obtaining the full flow of the machine penetration testing steps augmented as FARR Flow, we constructed a guidance reasoning test for the model by asking it what to do next when we have only the Findings. FARR Flow consists of multiple sets of Findings, Actions, Reasonings, and Results, as shown in Figure 10, where F is Finding, A is Action, R is Reasoning, and the last R is Result.
{adjustwidth}
[…]
Figure 10: FARR Flow evaluation.
@@
Figure  10 also illustrates how we evaluate the model using FARR Flow. At each step, we utilize the Findings as known information. We formulate questions for the model to predict the most likely action to do based on the current hint, which is constructed from the accumulated previous and current findings. The model’s suggested response is then evaluated against Llama 3 70B to assess the relevance of the answer across three aspects: whether the model’s answer anticipates the same outcome, targets the same service, and focuses on the same vulnerability type as the reference. Each criterion is rated with up to three points for relevance. Using these three aspects, we measure the average performance of each model on each Hack The Box machine in the FARR Flow evaluation.
@@
Our guidance reasoning test using the FARR Flow evaluation algorithm is shown in Algorithm 1. At the first question, the user does not know about the target machine, often only the IP information or open ports. Therefore, we only provide the current findings. However, regardless of the suggested answer to the following question, we will still provide the model with the result of the previous findings to ensure the model understands the current situation. Section 5.3 discusses the possibility of improvement in this question formulation.
@@
     current_findings $\leftarrow$ flow[‘‘Findings’’]
     current_result $\leftarrow$ flow[‘‘Result’’]
@@
CIPHER training data use the 0xdf write-up 0xd dataset until early September 2023. We took whole machine write-up samples from mid-September 2023 until the FARR Flow benchmark development started in early May 2024 without cherry-picking. We only excluded incomplete machine write-ups and separated them, resulting in 35 machine write-ups augmented as FARR Flow for the reasoning test, with different difficulties as shown in Figure 11a. Note that these machines are completely unknown to CIPHER. Overall, there are 35 Hack The Box machines with 136 vulnerability topics, all in 2124 formulated questions.
@@
Figure 11b presents different attack vectors in this evaluation. The scope includes diverse vulnerabilities and techniques used to penetrate the machine. Each machine has a different attack vector to achieve a foothold from easy to insane level. Various privilege escalation techniques are needed to achieve better scores, and CVE knowledge is necessary to access some machines. This covers better coverage and penetration testing dynamics than traditional question-and-answer evaluation.
{adjustwidth}
[…]
@@
Figure 11: Hack The Box machines: (a) Hack The Box machines selected for challenges sorted by difficulties, (b) Hack The Box machines selected for evaluation and corresponding attack vector scope.
@@
Figure 12 shows real example of a question augmented from FARR Flow from a Hack The Box Keeper machine, where it contains a list of findings and results to provide a situation of the user’s current penetration testing progress. Based on the question, the last situation is the current findings in the current FARR step; the model needs to understand the current progress first and then suggest the most likely action to be taken next. We ask for the most likely action to extract the best action suggested from the model to measure the accuracy of the vulnerability points, the service targeted, and the expected outcome from the action.
The prompt question in Figure 12 provides valuable information about the current status of the penetration testing progress, including open ports, service versions, subdomains, running instances, credentials, and both failed and successful attempts. These details are typically noted by penetration testers. For beginners in penetration testing, however, this information can be overwhelming, making it difficult to determine the appropriate next steps. In real-world scenarios, beginners often turn to forums to ask experts or more experienced penetration testers for hints, or what is commonly referred to as a “nudge” in the community.
{adjustwidth}
[…]
Figure 12: Real FARR Flow augmented question from Hack The Box Keeper machine. Notes (***: the rest of middle part findings.)
This is where CIPHER bridges the gap. According to the write-ups augmented in FARR Step Figure 12, the current finding is “Persistent .NET version issues on Linux,” and the correct action is to “Use Docker to create a consistent development environment with the correct .NET version.” However, all three of the other models failed to suggest this. Llama 3 70B, for instance, suggested attempting to extract a password, which is irrelevant. GPT-3.5 Turbo proposed analyzing a memory dump, which does not make sense since it was already stated that the tool could not run. Qwen 1.5 72B suggested installing .NET on Linux, which is also incorrect because the question clearly states that there is a persistent issue with .NET on Linux. CIPHER, however, responded correctly and aligned with the reference by suggesting that the user run the tool in Docker with the correct .NET tools. When provided with complex information from the penetration testing process, CIPHER was able to accurately determine the next step, outperforming even models with ten times the number of parameters, such as Llama 3 70B, GPT-3.5 Turbo, and Qwen 1.5 72B.
@@
The most relevant models are selected for this evaluation, including available small pentesting or cybersecurity models, and we also included the best open-source state-of-the-art (SOTA) model as the judge: Llama 3 70B. To ensure our benchmark is compatible with larger parameter models, we included the Llama 3 70B and Qwen 1.5 72B Chat scores in our benchmark. This allows us to assess how well our model’s guidance compares to larger models out of CIPHER’s size.
@@ 4.3.1 Penetration Testing Guidance Accuracy Performance
Table 9 shows the three criteria measured in our FARR Flow evaluation for each model tested. Our model CIPHER achieves the highest average score compared to other penetration testing models and large SOTA general models like Llama 3 70b and Qwen1.5 72B. Llama 3 70B scores slightly higher (by 0.19) in outcome prediction, but CIPHER guides better in service aimed and vulnerability targeted, with a 0.52 higher score overall at 52.59. However, compared to other penetration testing models similar to CIPHER’s size, there is a significant increase in performance, as big as 9.75% from Colibri 0.1 8B.
@@
Table 9: Model performance on FARR Flow reasoning test.
@@
whiterabbitneo-7b Whi 25.81 38.78 27.49 30.69
lily-7b seg 29.35 43.77 35.57 36.23
qwen-7b Qwe 36.50 51.18 40.11 42.60
colibri-0.1-8b Cyb 40.85 56.61 46.30 47.92
hermes2pro-llama3-8b Nou 41.27 58.13 47.15 48.85
mistral-ins-0.2-7b Mis 41.36 59.91 46.60 49.29
qwen-72b Qwe 43.57 59.05 46.84 49.82
gpt-3.5-turbo Ouyang et al. (2022) 43.57 59.31 47.45 50.11
llama-3-70b-ins Met 44.92 61.67 49.62 52.07
CIPHER DPO 7B (Ours) 44.73 62.65 50.39 52.59
@@
In summary, most penetration testing models struggle to handle the large volume of information and often suggest inaccurate next steps. However, state-of-the-art models like Llama 3 (70B) achieve the highest scores in outcome prediction. Despite this, our CIPHER DPO (7B) model outperforms even the largest models in terms of suggestion accuracy. This means that CIPHER guides users more effectively, enabling them to penetrate vulnerable machines faster, achieve a foothold, and gain the highest privilege accounts more efficiently than models ten times its size.